OpenStudy (anonymous):

Linda joined cardboard pieces to make a square pyramid, as shown below. If the dimensions of the pyramid are made four times greater, which statement is true about the surface area of the new pyramid?

OpenStudy (anonymous):

A square pyramid is shown with length of the square base sides equal to 9 inches and the altitude of one of the lateral triangular faces as 6 inches.

OpenStudy (anonymous):

waiting for the pic...

OpenStudy (anonymous):

Anything?

OpenStudy (anonymous):

usually it takes a while like the previous ones.. but still nothing...:(

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I might have an idea.

OpenStudy (anonymous):

Volume of a pyramid=1/3bh

OpenStudy (anonymous):

The base is 9 while the slant height is 6

OpenStudy (anonymous):

54/3=18

OpenStudy (anonymous):

If the dimensions were multiplied by 4, then the base would be 36 and the height would be 24

OpenStudy (anonymous):

36 times 24=864/3=288

OpenStudy (anonymous):

288/18=16!!!!!!!!!!

OpenStudy (anonymous):

It will be sixteen times the original pyramid’s surface area.

OpenStudy (anonymous):

I got this one, sir.